# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ Appendix I Models We Use in Experiments
choice2
choice3
choice4
relative_part
confuse rate
pp_part
confuse rate
compound_part
confuse rate
adverb_part
confuse rate
3-4B Llama Llama-3.2-3B Baseline zeroshot 0.537 68.59 23.05 8.36 25.26 35.52 56.68 33.33
fewshot 0.387 78.35 19.33 2.32 23.55 23.79 48.01 31.43
After SFT zeroshot 0.208 82.21 14.90 2.88 11.26 17.24 27.08 12.38
Llama-3.2-3B-Instruct Baseline zeroshot 0.510 67.51 30.14 2.35 29.01 33.45 47.29 30.48
fewshot 0.379 70.79 28.42 0.79 26.62 22.07 33.21 33.33
After SFT zeroshot 0.192 84.38 12.5 3.12 9.56 12.76 29.96 13.33
Qwen Qwen2.5-3B Baseline zeroshot 0.501 66.14 30.08 3.78 31.40 33.45 40.07 30.48
fewshot 0.392 71.76 23.16 5.09 25.26 22.76 38.63 33.33
After SFT zeroshot 0.254 84.65 14.17 1.18 14.68 23.45 33.57 10.48
Qwen2.5-3B-Instruct Baseline zeroshot 0.455 60.75 37.72 1.54 25.60 32.07 28.16 29.52
fewshot 0.241 75.10 24.48 0.41 16.04 14.83 24.55 21.90
After SFT zeroshot 0.236 80.93 16.95 2.12 12.97 21.03 29.24 10.48
7-8B Llama Llama-3.1-8B Baseline zeroshot 0.523 70.80 22.90 6.30 29.35 34.83 53.43 34.29
fewshot 0.357 81.01 16.20 2.79 22.53 21.38 47.29 29.52
After SFT zeroshot 0.149 80.54 16.11 3.36 9.56 10.34 18.77 9.52
Llama-3.1-8B-Instruct Baseline zeroshot 0.436 67.05 32.04 0.92 29.01 30.69 30.69 32.38
fewshot 0.297 77.18 22.15 0.67 22.53 19.31 28.88 26.67
After SFT zeroshot 0.145 84.83 13.10 2.07 9.22 11.72 18.05 11.43
gemma gemma-7b Baseline zeroshot 0.508 68.17 28.88 2.95 28.67 33.10 46.93 35.24
fewshot 0.382 83.81 14.62 1.57 26.28 23.45 49.10 38.10
After SFT zeroshot 0.153 77.78 20.26 1.96 9.56 10.34 19.49 6.67
gemma-7b-it Baseline zeroshot 0.412 61.99 34.14 3.87 17.06 30.00 34.66 21.90
fewshot 0.298 65.22 32.44 2.34 14.33 16.55 29.96 20.95
After SFT zeroshot 0.162 77.16 20.37 2.47 10.24 11.72 18.77 8.57
Qwen Qwen2.5-7B Baseline zeroshot 0.501 66.14 30.08 3.78 31.40 33.45 40.07 30.48
fewshot 0.387 74.23 24.23 1.55 24.91 24.48 41.52 27.62
After SFT zeroshot 0.180 85.56 13.33 1.11 14.33 15.52 20.22 10.48
Qwen2.5-7B-Instruct Baseline zeroshot 0.333 57.19 40.42 2.40 16.38 22.07 20.58 20.95
fewshot 0.299 68.67 31.33 0.00 19.45 20.00 25.99 18.10
After SFT zeroshot 0.217 84.79 14.29 0.92 14.68 15.86 28.88 14.29
Mistral Mistral-7B-v0.3 Baseline zeroshot 0.521 72.03 22.61 5.36 26.96 34.83 58.12 33.33
fewshot 0.372 79.09 18.23 2.68 22.87 22.76 48.01 27.62
After SFT zeroshot 0.161 83.85 13.66 2.48 8.87 11.03 25.27 6.67
Mistral-7B-Instruct-v0.3 Baseline zeroshot 0.333 57.49 41.92 0.60 20.48 21.72 18.05 18.10
fewshot 0.236 70.34 29.24 0.42 13.65 14.14 25.63 13.33
After SFT zeroshot 0.185 83.78 14.59 1.62 9.56 12.41 29.24 9.52
Table 17: Incorrect choice distribution and confusion analysis in Thunder-NUBench across 3-4B/7-8B, pretrained/instruction-tuned models.

### v3

@@ 8 Paraphrased sentence:
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2b baseline zero-shot 0.563 74.08 21.41 4.51 29.17 32.14 67.69 44.19
1-shot 0.558 74.54 21.34 4.13 32.37 31.49 63.27 45.16
5-shot 0.474 81.61 15.89 2.51 26.28 30.19 55.78 48.06
10-shot 0.421 82.49 16.01 1.51 23.72 23.38 48.64 48.06
sft zero-shot 0.268 84.91 14.20 0.89 13.14 18.51 26.87 35.48
gemma-1.1- 2b-it baseline zero-shot 0.612 59.20 32.25 8.55 31.09 28.25 55.44 35.48
1-shot 0.565 66.76 28.47 4.77 32.37 29.55 55.10 39.35
5-shot 0.471 66.50 27.44 6.06 25.64 24.35 47.96 31.94
10-shot 0.468 68.98 26.44 4.58 26.28 22.40 49.66 35.48
sft zero-shot 0.250 81.27 17.78 0.95 11.86 15.26 26.87 30.00
gemma-7b baseline zero-shot 0.546 72.28 24.24 3.48 29.49 32.14 59.18 42.90
1-shot 0.535 77.74 18.10 4.15 30.77 31.82 57.82 51.61
5-shot 0.437 84.39 13.79 1.81 25.64 26.95 47.96 51.94
10-shot 0.369 84.95 13.33 1.72 22.44 23.05 35.71 48.06
sft zero-shot 0.203 81.64 16.80 1.56 10.26 11.36 18.03 28.71
gemma-1.1- 7b-it baseline zero-shot 0.337 55.06 38.35 6.59 9.29 14.29 26.53 26.77
1-shot 0.310 69.57 26.09 4.35 14.10 13.31 26.53 35.16
5-shot 0.279 73.01 25.00 1.99 13.14 13.64 20.07 37.10
10-shot 0.232 73.72 24.91 1.37 10.58 7.79 15.31 36.77
sft zero-shot 0.214 80.74 17.78 1.48 10.90 11.04 17.35 31.94
Table 15: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2b baseline zero-shot 0.803 274 33.29 33.42 33.29 20.51 23.05 21.43 15.48
1-shot 0.750 0 31.18 29.6 39.22 22.44 26.62 23.47 23.87
5-shot 0.722 0 30.63 23.82 45.55 20.19 23.70 28.23 19.35
10-shot 0.736 0 29.42 22.74 47.84 21.47 23.70 23.81 20.32
sft zero-shot 0.741 6 32.54 34.48 32.97 22.44 27.92 25.51 22.90
gemma-1.1- 2b-it baseline zero-shot 0.738 4 34.23 27.86 37.90 24.04 28.25 25.85 25.48
1-shot 0.654 0 46.67 18.91 34.42 29.81 35.71 30.61 29.68
5-shot 0.631 0 48.68 14.59 36.73 32.05 37.34 27.55 29.35
10-shot 0.660 0 43.87 11.06 45.07 29.49 34.42 25.85 29.35
sft zero-shot 0.994 1,239 85.71 0 14.29 0.96 0.97 1.02 0.97
gemma-7b baseline zero-shot 0.612 195 57.89 9.88 32.24 17.31 19.81 44.90 28.06
1-shot 0.351 0 69.07 12.19 18.74 20.51 14.61 40.82 24.84
5-shot 0.307 0 77.52 15.50 6.98 22.76 17.86 35.03 22.90
10-shot 0.288 0 81.54 13.50 4.96 24.36 16.23 31.29 25.16
sft zero-shot 0.278 1 70.57 12.57 16.86 11.54 14.29 28.91 26.45
gemma-1.1- 7b-it baseline zero-shot 0.355 0 64.51 20.09 15.40 16.99 15.26 40.14 22.90
1-shot 0.307 0 63.05 15.76 21.19 16.03 11.36 34.01 19.03
5-shot 0.283 0 63.87 18.77 17.37 16.99 9.42 29.59 19.03
10-shot 0.289 0 65.93 17.31 16.76 19.23 8.77 30.61 20.32
sft zero-shot 0.269 2 76.26 16.91 6.820 22.12 14.61 27.21 20.32
Table 16: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.2- 3B baseline zero-shot 0.556 70.04 23.11 6.85 24.68 30.19 62.59 44.19
1-shot 0.519 75.99 20.49 3.52 25.32 30.19 60.54 47.42
5-shot 0.419 78.41 18.94 2.65 20.19 25.32 46.60 43.87
10-shot 0.401 80.20 16.83 2.97 22.44 24.03 43.20 43.23
sft zero-shot 0.267 82.79 15.43 1.78 14.42 17.53 25.85 33.55
Llama-3.2- 3B-Instruct baseline zero-shot 0.528 72.67 23.87 3.45 23.72 29.55 50.68 54.84
1-shot 0.502 75.36 23.22 1.42 28.85 27.27 48.98 51.29
5-shot 0.441 76.08 22.84 1.08 25.64 26.62 39.80 46.45
10-shot 0.410 76.79 21.86 1.35 25.64 24.68 35.03 44.52
sft zero-shot 0.255 83.80 14.95 1.25 12.82 13.31 27.55 34.52
Llama-3.1- 8B baseline zero-shot 0.562 70.62 21.33 8.05 25.64 30.84 64.29 43.87
1-shot 0.507 77.78 18.15 4.07 25.96 29.87 59.18 48.39
5-shot 0.393 83.87 14.31 1.81 20.83 22.40 47.62 45.81
10-shot 0.358 84.51 13.72 1.77 20.19 23.05 40.82 41.29
sft zero-shot 0.203 85.55 12.50 1.95 8.97 11.69 18.71 32.26
Llama-3.1- 8B-Instruct baseline zero-shot 0.488 70.57 26.18 3.25 24.36 25.97 54.76 37.74
1-shot 0.419 73.86 24.43 1.70 22.76 25.97 44.56 34.84
5-shot 0.347 80.55 18.76 0.69 18.91 19.16 39.80 37.74
10-shot 0.329 81.20 18.31 0.48 19.55 20.45 34.01 36.45
sft zero-shot 0.244 86.04 12.34 1.62 12.18 12.66 28.23 33.87
Table 17: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.2- 3B baseline zero-shot 0.688 0 42.21 25.95 31.83 22.12 33.12 32.65 31.94
1-shot 0.570 0 62.59 14.74 22.67 31.09 32.14 45.58 38.71
5-shot 0.548 0 74.10 14.33 11.58 34.62 35.06 50.34 47.74
10-shot 0.500 0 78.13 13.00 8.87 35.90 31.82 43.88 49.68
sft zero-shot 0.658 0 39.40 33.01 27.59 23.40 31.17 27.89 24.52
Llama-3.2- 3B-Instruct baseline zero-shot 0.474 8 63.90 10.34 25.76 19.55 28.25 60.20 16.77
1-shot 0.461 0 71.60 12.56 15.83 23.40 25.97 61.22 26.77
5-shot 0.472 0 74.62 10.92 14.45 29.81 28.25 63.27 25.16
10-shot 0.467 0 73.85 10.36 15.79 33.01 25.00 62.24 23.23
sft zero-shot 0.437 87 61.85 7.11 31.03 11.22 18.51 43.54 21.61
Llama-3.1- 8B baseline zero-shot 0.541 0 47.95 6.74 45.31 23.08 24.03 41.50 19.03
1-shot 0.454 0 67.66 4.55 27.80 27.56 28.25 47.62 23.87
5-shot 0.362 0 78.95 6.58 14.47 30.45 21.75 40.82 25.16
10-shot 0.316 0 82.46 6.02 11.53 30.77 21.75 31.97 23.23
sft zero-shot 0.444 48 58.79 27.34 13.87 20.83 20.13 37.41 20.65
Llama-3.1-8B- Instruct baseline zero-shot 0.256 12 76.53 15.11 8.36 16.67 18.83 19.05 23.23
1-shot 0.290 0 71.51 21.37 7.12 19.23 18.18 29.25 19.03
5-shot 0.269 0 74.04 21.83 4.13 20.83 16.88 25.17 19.35
10-shot 0.290 0 80.00 14.52 5.48 25.00 20.78 26.87 22.90
sft zero-shot 0.254 31 72.32 19.03 8.65 15.06 17.86 21.09 14.52
Table 18: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B- v0.3 baseline zero-shot 0.575 76.97 18.48 4.55 26.60 32.14 66.33 58.39
1-shot 0.532 78.39 17.59 4.02 26.92 33.12 59.52 53.23
5-shot 0.439 83.57 14.98 1.44 22.12 25.65 51.70 52.58
10-shot 0.401 84.39 13.83 1.78 21.15 24.68 47.28 47.10
sft zero-shot 0.227 88.81 9.09 2.10 8.65 13.31 29.93 31.61
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.381 67.29 30.83 1.88 13.46 18.51 30.95 42.90
1-shot 0.365 72.61 26.09 1.30 15.38 20.45 34.69 39.03
5-shot 0.331 72.66 26.86 0.48 14.10 15.58 31.97 37.74
10-shot 0.304 73.63 25.85 0.52 12.18 14.29 30.61 35.48
sft zero-shot 0.236 85.52 13.80 0.67 10.26 12.01 27.21 33.87
Table 19: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B- v0.3 baseline zero-shot 0.624 351 53.21 21.56 25.23 16.99 18.18 26.53 14.52
1-shot 0.410 0 67.12 16.63 16.25 30.77 24.68 28.91 29.03
5-shot 0.306 0 83.16 11.92 4.92 26.28 24.35 24.49 29.68
10-shot 0.272 0 85.42 8.45 6.12 23.40 22.73 21.09 28.39
sft zero-shot 0.614 0 40.05 29.46 30.49 22.44 27.92 28.57 22.58
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.358 1 65.56 25.78 8.67 27.24 23.05 27.55 18.71
1-shot 0.370 0 61.03 34.90 4.07 27.24 20.78 26.19 19.03
5-shot 0.335 0 68.96 29.86 1.18 28.85 19.81 26.19 20.32
10-shot 0.331 0 71.22 27.10 1.68 28.53 22.08 25.51 20.97
sft zero-shot 0.351 1 67.57 21.32 11.11 23.08 23.05 33.67 18.06
Table 20: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen2.5- 3B baseline zero-shot 0.542 67.69 26.02 6.29 24.68 30.19 55.10 42.26
1-shot 0.522 67.63 25.53 6.84 24.36 28.25 55.78 38.06
5-shot 0.481 70.84 23.89 5.27 25.00 28.57 50.00 37.74
10-shot 0.437 73.14 22.14 4.72 23.40 25.32 45.58 38.06
sft zero-shot 0.276 85.34 13.22 1.44 14.74 20.78 30.95 30.97
Qwen2.5- 3B-Instruct baseline zero-shot 0.466 68.03 28.57 3.40 19.23 28.25 42.86 40.97
1-shot 0.393 67.68 28.28 4.04 16.67 20.45 42.52 30.65
5-shot 0.355 77.01 21.88 1.12 19.55 22.40 39.46 31.94
10-shot 0.351 77.83 21.04 1.13 21.15 23.05 35.71 32.90
sft zero-shot 0.300 85.19 13.49 1.32 15.71 22.40 33.33 34.19
Qwen2.5- 7B baseline zero-shot 0.520 70.73 24.54 4.73 24.68 31.17 53.06 43.55
1-shot 0.534 71.47 25.56 2.97 29.17 31.17 56.12 41.61
5-shot 0.492 71.13 25.81 3.06 24.36 28.25 51.70 40.65
10-shot 0.447 74.07 23.98 1.95 24.04 24.68 45.58 42.58
sft zero-shot 0.240 87.75 11.59 0.66 15.06 15.91 22.45 33.23
Qwen2.5- 7B-Instruct baseline zero-shot 0.458 65.16 31.54 3.29 18.27 29.22 39.80 36.13
1-shot 0.438 69.02 30.07 0.91 22.12 23.70 43.88 35.48
5-shot 0.378 71.85 27.94 0.21 19.55 19.16 37.41 36.13
10-shot 0.335 73.76 25.77 0.47 19.23 18.18 30.61 34.19
sft zero-shot 0.282 88.76 10.96 0.28 16.03 19.48 31.29 36.77
Table 21: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen2.5- 3B baseline zero-shot 0.500 0 38.99 19.49 41.52 23.08 19.16 26.87 11.61
1-shot 0.463 0 45.89 27.23 26.88 25.00 21.43 26.53 14.84
5-shot 0.431 0 44.75 25.60 29.65 24.04 19.81 23.13 12.58
10-shot 0.439 0 45.85 23.10 31.05 25.00 21.10 25.17 11.94
sft zero-shot 0.373 2 51.07 13.25 35.68 17.31 15.58 26.19 19.35
Qwen2.5- 3B-Instruct baseline zero-shot 0.380 0 49.27 40.29 10.44 20.19 15.91 24.15 17.10
1-shot 0.419 0 53.22 32.39 14.39 26.28 19.16 28.23 18.39
5-shot 0.398 2 54.40 31.60 14.00 27.24 19.16 29.25 13.55
10-shot 0.388 0 56.44 29.24 14.31 26.60 20.13 30.61 13.23
sft zero-shot 0.274 1 59.13 34.49 6.38 12.18 13.31 21.43 20.00
Qwen2.5- 7B baseline zero-shot 0.377 0 68.21 15.79 16.00 28.53 22.73 36.39 18.71
1-shot 0.392 0 70.65 15.59 13.77 33.97 26.30 38.10 16.13
5-shot 0.377 0 72.00 15.16 12.84 33.65 23.05 38.78 16.77
10-shot 0.372 0 73.13 14.93 11.94 32.69 23.05 40.48 16.45
sft zero-shot 0.344 0 82.03 13.13 4.84 30.77 24.35 43.54 18.39
Qwen2.5-7B- Instruct baseline zero-shot 0.324 0 62.84 29.34 7.82 23.72 22.73 23.47 14.19
1-shot 0.363 0 61.79 31.66 6.55 30.45 21.10 27.21 13.87
5-shot 0.368 0 66.38 26.51 7.11 31.41 19.81 34.69 15.16
10-shot 0.390 0 67.07 25.81 7.11 34.62 20.78 35.03 17.74
sft zero-shot 0.207 0 78.54 17.62 3.83 18.59 20.13 17.69 10.65
Table 22: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gpt-4o- mini baseline zero-shot 0.357 0 34.67 65.33 0 17.31 8.44 13.27 11.94
1-shot 0.317 0 40.00 59.75 0.25 19.87 7.47 13.95 10.97
5-shot 0.316 0 38.35 61.15 0.50 19.23 7.14 11.90 11.61
10-shot 0.298 0 42.02 57.98 0 19.23 8.12 12.93 11.29
gpt-4.1- mini baseline zero-shot 0.203 0 51.95 48.05 0 14.10 6.49 8.16 14.52
1-shot 0.184 0 64.66 35.34 0 16.35 6.17 10.20 16.13
5-shot 0.156 0 65.48 34.52 0 13.46 6.17 7.14 15.16
10-shot 0.153 0 64.77 35.23 0 14.74 5.19 4.76 15.81
claude-3-5- haiku-latest baseline zero-shot 0.278 0 67.81 30.48 1.71 22.76 14.29 15.99 24.52
1-shot 0.477 434 83.93 15.48 0.60 20.19 6.49 10.54 8.71
5-shot 0.527 574 95.56 4.44 0 13.78 3.57 5.44 5.16
10-shot 0.368 324 94.29 5.00 0.71 20.83 6.17 5.44 10.32
Table 23: Error rates, incorrect choice distributions, and local negation confusion rates for the API models under zero-shot and few-shot conditions, evaluated in the option-selection setting.